OpenStudy (anonymous):

sharon is the yougest in the family anne is 3 years older than sharon and tommy is 5 years older than anneall toghther they equal 62 how old is sharon

OpenStudy (radar):

Welcome to Open Study. Sometimes those written problems can be confusing. But, if you label what you are told, and some algebra, you can solve.

OpenStudy (radar):

In this problem we are told that Sharon is the youngest so lets label Sharon's age as x. Let x = Sharon's age. we are not told that anne is 3 years older than Sharon Let x + 3 equal = Anne's age. Tommy is 5 years older than Anne Let x + 3 + 5 = Tommy's age. They sum up to 62 years. This is all we need we now add their ages: x + x + 3 + x + 3 +5 I will enclose each age in parenthese so you can see it better (x) + (x + 3) + (x + 3 + 5) = 62 Now it is just algebra which you have already done. 3x + 11 = 62 3x = 62 - 11 3x = 51 x = 17 (Sharon's age) x + 3 = 20 (Anne's age) and x + 3 + 5 = 25 (Tommy's age. Do they add up to 62? 17 + 20 +25 = 62 Yes is Anne 3 years older than sharon? is tommy 5 year older than anne?